OpenStudy (maheshmeghwal9):

Prove that the polynomial \[x^{3n}+x^{3m+1}+x^{3k+2}\] is exactly divisible by \[x^2+x+1\] if m, n, k are any non-negative integers. I know only this much that \[x^2+x+1\] becomes zero at \[\omega \space and \space \omega ^2\] .

OpenStudy (maheshmeghwal9):

@waterineyes Plz help:)

OpenStudy (anonymous):

Frankly speaking I have no clue how to do this but I am thinking of it..

OpenStudy (maheshmeghwal9):

ok np :) but i have done something

OpenStudy (anonymous):

Yes you can show that to me..

OpenStudy (anonymous):

Can anyone here help us...

OpenStudy (anonymous):

we see that by substuting omega in the first equation the result is zer0......similarly with omega squared its the same result.......got any idea???

OpenStudy (maheshmeghwal9):

I have done put omega in \[1+x+x^2\]\[1+ \omega + \omega ^2=0\]so put x = omega in \[\omega^{3n}+\omega^{3m+1}+\omega^{3k+2}=1+\omega +\omega^2=0\]

OpenStudy (anonymous):

Now try with w^2 as Prophet is saying..

OpenStudy (anonymous):

yup.....we get zero even by substituting omega squared........ that means .....by factor theorem first equation can be written as \[(x-\omega)(x-\omega^{2})(some polymonial)\]

OpenStudy (maheshmeghwal9):

i m getting same with omega squred

OpenStudy (anonymous):

and second equation is \[(x-\omega)(x-\omega^{2})\]

OpenStudy (maheshmeghwal9):

but we have to prove:)

OpenStudy (maheshmeghwal9):

how can we prove tht statement?

OpenStudy (anonymous):

divide them both......by cancelling \[(x-\omega)(x-\omega^{2})\] we get some polymonial ...... hence proved!!!!

OpenStudy (maheshmeghwal9):

so both equation can be written as \[(x- \omega)(x-\omega^2)\]therefore \[\frac {\cancel{(x-\omega)(x-\omega^2)}}{\cancel{(x-\omega)(x-\omega^2)}}=1.\]hence proved. Am i right?

OpenStudy (anonymous):

no!!!!! the first equation can be written as (x-w)(x-w^2)*(some polynomial)........... its same as in wht is 24 divisible by 3....... its because 3 is a factor of 24..... 24=3*12 \[\frac{24}{3}=\frac{3*12}{3}\]

OpenStudy (maheshmeghwal9):

but the first eq. is also like this \[1+\omega+\omega^2\]then it must be only \[(x-\omega)(x-\omega^2).\]

OpenStudy (anonymous):

by substituting and getting zero we mean that omea and omega squared are factors of the first equation......just google 'factor theorem' .....u'll know what i mean...

OpenStudy (maheshmeghwal9):

ohh i seeeeeeeeeee thanx a lot now i gt wt u r saying:)

OpenStudy (anonymous):

pleasure!!!